# Analysis Theory of Data Economy : Dataization, Technological Progress and Dynamic General Equilibrium

### v1

@@ 4 Empirical Evidence and Analysis
From the numerical simulation analysis and the results of Figure 2, it can be seen that an increase in the proportion of total output datatization $\theta$ can promote technological progress, i.e., an increase in $\eta$, when the equilibrium capital stock $k^{*}$ and the equilibrium consumption $c^{*}$ are at the steady-state level, i.e., at fixed values. Accordingly, we present the core proposition of the empirical analysis section:
@@
This paper uses the Staggered Difference-in-Difference model to explore how Digital and Tech change after the implementation of open data policies. Using the open data pilot cities as the core experimental subjects, the product of the policy implementation time dummy variable and the experimental group dummy variable constitutes the DID. The time dummy variable is: if a provincial region has conducted an open data pilot in a certain year, the region will be assigned a value of 1 for this year and the following years, otherwise it will be 0. The experimental group dummy variable is: if a provincial region has conducted an open data pilot in a certain year, the region will be assigned a value of 1 from this year onwards, and prior to this year the region is assigned a value of 0, which is the control group. In this paper, two macro policies related to data openness are selected, the first one is “Establishment of Regional Government Data Open Platform” (DID1), and the second one is “Establishment of Regional Comprehensive Pilot Zones for Cross-border E-commerce” (DID2). The time cut-off points for the implementation of both policies are taken from public documents on the official websites of the central and local governments.
@@
The results of the benchmark experiment [^5] Standard errors in parentheses: ${}^{*}p<0.05$,${}^{**}p<0.01$,${}^{***}p<0.001$ are shown in Table 1, and the results of the policy effects experiment are shown in Figure 5, Figure 6, Figure 7, Figure 8, they show the shocks of datatization policy on technical progress variables under two-way fixed effects. From them, we find that datatization policies can significantly promote technological progress, consistent with the core proposition of the article:
Columns (1)-(3) show the impact of the first type of datatization policy, i.e., “government data platform establishment”, on technological progress: it is found that the first type of datatization policy has a significant effect on the development of the digital economy and the number of patents granted in the digital economy. Although the promotion effect of the first type of policy on the development of digital economy is no longer significant after adding the time fixed effect, we can know through Figure 5 that this is because there is a lag in the promotion effect of the first type of policy on the development of digital economy, which may be related to the implementation of the policy of data openness, the establishment of databases and data management platforms need time. After the 4th period of the first type of policy shock, the effects of such policies on the development of digital economy are all significantly positive. Figure 6 presents the impact effect of the first type of policy on the number of patents granted in the digital economy. The coefficients are not significant before the policy shock, and after the policy shock, the coefficients are all significantly positive, which fully demonstrates that the first type of datatization policy can effectively promote the enhancement of patent granting volume of digital economy.
Columns (4)-(6) show the impact of the second type of datatization policy, i.e., the “establishment of comprehensive cross-border e-commerce pilot zones”, on technological progress: from which it is found that the second type of datatization policy has, in the case of two-way fixed effects, a significant boosting effect. As can be seen from Figure 7, the promotion effect of the second type of policy on digital economy development is similar to the case of the first type of policy, with a lag period, and it is only after the 4th period of the second type of policy shock that it is able to significantly promote the development of the digital economy, and it is therefore reasonable to assume that this is due to the time needed for the establishment of databases and data management platforms. As can be seen from Figure 8, the second type of policy has a very good positive shock effect on the number of patents granted in the digital economy, so it is reasonable to believe that the second type of policy can significantly increase the number of patents granted in the digital economy.
@@ 5 Conclusion
In summary, the present paper identifies three analysis theories for the data economy. Firstly, technological advances have the potential to significantly promote the equilibrium capital stock of firms. That is, the increase of $\eta$ can lead to an increase in $k^{*}$. However, the degree of data openness has been demonstrated to be inversely correlated with equilibrium capital stock, that is, the increase of $\theta$ will lead to a decrease in $k^{*}$, which means the greater the degree of data openness, the less the equilibrium capital stock, and the stronger the substitution effect of data on capital factors. Secondly, for the equilibrium consumption of households, the impact of technological progress on equilibrium consumption is a typical U-shaped curve, and the policy regarding data openness will make it more persistent, gradually transforming it from U-shaped curve to S-shaped curve. That means the increase of data openness $\theta$ could assist in enhancing the threshold of technological progress $\eta$ for equilibrium consumption $c^{*}$ growth. Finally, we generalize the nonlinear effects of data openness policy on the impact of technological progress on dynamic general equilibrium to an important theory of data economy: data openness could profoundly promote the progress of digital technology and the development of digital economy. The proposition is formulated through numerical simulations, and four DID experiments are designed to demonstrate the core ideas. The empirical results from our experiments are consistent and significant with the theoretical analysis.

### v2

@@ 4 Empirical Evidence and Analysis
From the numerical simulation analysis and the results of Figure 2, it can be seen that an increase in the proportion of total output dataization $\theta$ can promote technological progress, i.e., an increase in $\eta$, when the equilibrium capital stock $k^{*}$ and the equilibrium consumption $c^{*}$ are at the steady-state level, i.e., at fixed values. Accordingly, we present the core proposition of the empirical analysis section:
@@
This paper uses the Staggered Difference-in-Difference model to explore how Digital and Tech change after the implementation of open data policies. Using the open data pilot cities as the core experimental subjects, the product of the policy implementation time dummy variable and the experimental group dummy variable constitutes the DID. The time dummy variable is: if a provincial region has conducted an open data pilot in a certain year, the region will be assigned a value of 1 for this year and the following years, otherwise it will be 0. The experimental group dummy variable is: if a provincial region has conducted an open data pilot in a certain year, the region will be assigned a value of 1 from this year onward, and prior to this year the region is assigned a value of 0, which is the control group. In this paper, two macro policies related to data openness are selected, the first one is “Establishment of Regional Government Data Open Platform” (DID1), and the second one is “Establishment of Regional Comprehensive Pilot Zones for Cross-border E-commerce” (DID2). The time cut-off points for the implementation of both policies are taken from public documents on the official websites of the central and local governments.
@@
The results of the benchmark experiment [^5] Standard errors in parentheses: ${}^{*}p<0.05$,${}^{**}p<0.01$,${}^{***}p<0.001$ are shown in Table 1, and the results of the policy effects experiment are shown in Figure 5, Figure 6, Figure 7, Figure 8, they show the shocks of dataization policy on technical progress variables under two-way fixed effects. From them, we find that dataization policies can significantly promote technological progress, consistent with the core proposition of the article:
Columns (1)-(3) show the impact of the first type of dataization policy, i.e., “government data platform establishment”, on technological progress: it is found that the first type of dataization policy has a significant effect on the development of the digital economy and the number of patents granted in the digital economy. Although the promotion effect of the first type of policy on the development of digital economy is no longer significant after adding the time fixed effect, we can know through Figure 5 that this is because there is a lag in the promotion effect of the first type of policy on the development of digital economy, which may be related to the implementation of the policy of data openness, the establishment of databases and data management platforms need time. After the 4th period of the first type of policy shock, the effects of such policies on the development of digital economy are all significantly positive. Figure 6 presents the impact effect of the first type of policy on the number of patents granted in the digital economy. The coefficients are not significant before the policy shock, and after the policy shock, the coefficients are all significantly positive, which fully demonstrates that the first type of dataization policy can effectively promote the enhancement of patent granting volume of digital economy.
Columns (4)-(6) show the impact of the second type of dataization policy, i.e., the “establishment of comprehensive cross-border e-commerce pilot zones”, on technological progress: from which it is found that the second type of dataization policy has, in the case of two-way fixed effects, a significant boosting effect. As can be seen from Figure 7, the promotion effect of the second type of policy on digital economy development is similar to the case of the first type of policy, with a lag period, and it is only after the 4th period of the second type of policy shock that it is able to significantly promote the development of the digital economy, and it is therefore reasonable to assume that this is due to the time needed for the establishment of databases and data management platforms. As can be seen from Figure 8, the second type of policy has a very good positive shock effect on the number of patents granted in the digital economy, so it is reasonable to believe that the second type of policy can significantly increase the number of patents granted in the digital economy.
@@ 5 Conclusion
In summary, the present paper identifies three analysis theories for the data economy, we attempt to complement and refine the fundamental economic theory of productivity growth, technological progress, and dynamic general equilibrium in the field of data economics :
Firstly, technological advances have the potential to significantly promote the equilibrium capital stock of firms. That is, the increase of $\eta$ can lead to an increase in $k^{*}$. However, the degree of data openness has been demonstrated to be inversely correlated with equilibrium capital stock, that is, the increase of $\theta$ will lead to a decrease in $k^{*}$, which means the greater the degree of data openness, the less the equilibrium capital stock, and the stronger the substitution effect of data on capital factors. Secondly, for the equilibrium consumption of households, the impact of technological progress on equilibrium consumption is a typical U-shaped curve, and the policy regarding data openness will make it more persistent, gradually transforming it from U-shaped curve to S-shaped curve. That means the increase of data openness $\theta$ could assist in enhancing the threshold of technological progress $\eta$ for equilibrium consumption $c^{*}$ growth. Finally, we generalize the nonlinear effects of data openness policy on the impact of technological progress on dynamic general equilibrium to an important theory of data economy: data openness could profoundly promote the progress of digital technology and the development of digital economy. The proposition is formulated through numerical simulations, and four DID experiments are designed to demonstrate the core ideas. The empirical results from our experiments are consistent and significant with the theoretical analysis.